Find the inverse of the function
f(x) = 8/(x + 8)
Answer Choices:
1. x/(8+8x) 2. (-8x+8)/x 3. (8+8x)/x 4. not invertible
I keep getting -8+(8/x)

Answer by Alan3354(69443)

Find the inverse of the function
f(x) = 8/(x + 8)
---------
y = 8/(x+8)
==============
x = 8/(y+8) ----- solve for y
x*(y+8) = 8
y+8 = 8/x
y = (8/x) - 8
Your answer is correct, but it can be converted to #2
Y = (8-8x)/x
y = (-8x+8)/x